Solve the equation for n: nP2 = 6
Answer by ewatrrr(24785)

nP2 = n!/(n-2)! = 6
n(n-1) = 6
n^2 - n = 6
n^2 - n - 6 = 0
(n-3)(n+2) = 0
n = 3
3P2 = 6